Tsys increase nov,dec12
lbw Tsys (horcal) as well as the
16K stage dewar temperature (.ps) (.pdf):
the comparison of the sep12 and
dec12 calibration runs (.ps) (.pdf):
The system temperature for polA,polB in lbw
had been near 30k up to 17nov12. On 18nov12 the helium lines
became contaminated and caused the frig to stop. This caused the
pressure relief valve to open dumping room temperature air
into the dewar. After this, Tsys remained higher than normal.
lbw had 3 warmup cool down cycles:
- 18nov12 : when the valve opened
- 27nov12: warm up, cool down to try and fix the problem.
- 03dec12: when dewar was brought down to lab.
After reinstalling lbw on 07dec12, Tsys remained high:
- 17nov12:tsys: 30.6, 30.1 (POLA,POLB)
- 18nov12 lbw dewar pressure relief valve blew out.
- contamination in helium line caused refrig to lock up, then
- dewar warmed up. then it was cooled back down by
- 27nov12: dewar was warmed up. It was cooled back down by
- Tsys remained high after the cool down
- 04dec12: lbw system warmed up and brought down to the lab for
- 07dec12: lbw system reinstalled on telescope. started cooling
- 10dec12: system cool.
- 10dec12: x102 calibration runs show that Tsys is high. It is
not a problem with the cal values.
Plotting Tsys and dewar temps
The plots show lbw Tsys (horcal)
as well as the 16K stage dewar temperature (.ps) (.pdf):
After bringing lbw down to the lab, cleaning it,
and the reinstalling it, tsys remained higher than normal. This
could be a problem with the cal values, or a real increase in Tsys.
- Top: lbw Tsys using hcorcal. 01nov12 to 11dec12.
- Black is polA, red is polB
- Dashed lines show when various events occurred.
- You can see the stable 30K tsys for polA,B prior to 18nov12.
- middle: lbw Tsys blowup tsys 17nov12 - 11dec12.
- Bottom: 16k stage dewar temperature for 17nov12 -
- After the 1st cool down (19nov12), tsys remained about 35k
- polA then began to increase
- on 26nov12 the 16k temp jumped to about 50K.
- This was thought to be a monitor problem since the vapor
pressure gage showed the dewar to still be cold
- A problem with the power supply in the postamp chassis was
later found in the lab.
- Tsys also jumped at this time (which makes you think it was
a real increase in tsys.. unless the bad power supply was also
controlling the cal switching).
On 10dec12 an x102 calibration run was done using
lbw on the source B0518+165 (3C138). This data was then compared
with the an x102 run using B0518+165 done on 19sep12 (when tsys was
The plots show
the comparison of the sep12 and dec12 calibration runs (.ps) (.pdf):
- Page 1: gain,tsys,SEFD, average beam width for the to runs.
- red is sep12, black is dec12.
- The gains are about the same.
- Tsys and sefd are higher for the dec12 run.
- Page 2: coma, 1st sidelobe height, eta main beam, eta main
beam + first sidelobe
- Page 3: Pointing error vs az and za for the 2 runs.
- Page 4:Interpolate dec12 data to the sep12 positions.
- top: over plot SEFD vs za for the two dates.
- black is dec12, red is sep12.
- Bottom: plot the dec12/sep12 ratios vs za for the various
- Black: SEFD ratio
- Dec12 SEFD is about 16% higher than the sep12 values
- Red: Tsys ratio:
- Dec12 Tsys is about 18% higher than the sep12 values
- Green: Gain ratio:
- The dec12 gain is about 2% higher:
- Tsys increased after the nov18 pressure value opening
- PolA increased more than polB
- After reinstalling the dewar in the dome, Tsys remained higher
than the pre 18nov12 values:
- PolA remains higher than PolB
- comparing the SEFD's from sep12, dec12 shows:
- the SEFD increase is coming from an increase in Tsys and is
not a problem with the cal values changing.